# Kai Yao    Marc Juarez

### v1

@@ Abstract
Model fingerprint detection techniques have emerged as a promising approach for attributing AI-generated images to their source models, but their robustness under adversarial conditions remains largely unexplored. We present the first systematic security evaluation of these techniques, formalizing threat models that encompass both white- and black-box access and two attack goals: fingerprint removal, which erases identifying traces to evade attribution, and fingerprint forgery, which seeks to cause misattribution to a target model. We implement five attack strategies and evaluate 14 representative fingerprinting methods across RGB, frequency, and learned-feature domains on 12 state-of-the-art image generators. Our experiments reveal a pronounced gap between clean and adversarial performance. Removal attacks are highly effective, often achieving success rates above $80\%$ in white-box settings and over $50\%$ under constrained black-box access. While forgery is more challenging than removal, its success significantly varies across targeted models. We also identify a utility–robustness trade-off: methods with the highest attribution accuracy are often vulnerable to attacks. Although some techniques exhibit robustness in specific settings, none achieves high robustness and accuracy across all evaluated threat models. These findings highlight the need for techniques balancing robustness and accuracy, and identify the most promising approaches for advancing this goal.
@@ I Introduction
The widespread adoption of generative AI models for image generation has created an urgent need for content provenance mechanisms. These models—now integrated into creative tools [adobe2023firefly], content platforms [katsamakas2024gcp], and commercial APIs [openai2023dalle3]—can produce images that are almost visually indistinguishable from authentic photographs. While these technological advances enable new forms of creativity, they also come with risks for misinformation, fraud, and abuse, particularly when AI-generated content is passed as genuine [bird2023typology]. The challenge of reliably identifying and tracing the generated images has thus become critical for maintaining trust and accountability in digital media.
@@
Existing MFD techniques evaluated under various settings achieve over 90% accuracy for both deepfake detection [Marra19a, Nataraj19, Durall20, Dzanic20, Wang20, Corvi23] and model attribution [Marra18, Marra19b, Yu19, Wang20, Guarnera20, Girish21]. If these accuracies were to hold in realistic scenarios, they would prove useful in bridging the accountability gap in current deployments of generative models. Indeed, being able to identify deepfakes would help mitigate the risks of misinformation and scams, while attributing a harmful output to the specific model that generated it would make it possible to flag issues, understand their causes, and prevent similar problems in the future.
@@
We focus on model attribution rather than deepfake detection because it is a harder problem: model attribution does not merely require detecting the presence of model-induced patterns, but to detect architecture-specific differences across these patterns. Therefore, model attribution is at least as hard as deepfake detection and by evaluating MFD techniques in this setting, we evaluate the worst-case scenario for the fingerprinter, allowing us to draw conclusions for the other scenario: if the techniques prove robust against attacks on model attribution, they should remain robust for the simpler task of deepfake detection.
@@ II-A Model Fingerprint Detection
Model fingerprint detection (MFD) has gained significant attention as a passive approach to identifying images produced by generative models, such as GANs, VAEs, and diffusion models. Unlike watermarking, which embeds signals during generation, MFD relies on detecting inherent artifacts or “fingerprints” left by the model’s processes. These fingerprints can be detected in statistical irregularities of color and frequency patterns, as well as in feature representations.
@@
Frequency Domain. Generative models often introduce spectral anomalies, such as deficiencies in high-frequency components or unnatural power distributions, which frequency-based methods target through transformations like FFT or DCT. For example, approaches using azimuthally averaged power spectral densities from 2D FFTs have demonstrated near-perfect discrimination (close to 100% accuracy) for deepfake detection between GAN outputs and real images on CelebA and LSUN, detecting patterns in high-frequency components for models like DCGAN and CycleGAN [Durall20]. Other methods extract features from Fourier modes, such as decay exponents and magnitudes, achieving high accuracy (up to 99%) for deepfake detection on ProGAN-generated faces [Dzanic20]. Giudice et al. instead rely on DCT analysis, extracting a 63-dimensional fingerprint vector from the means of AC coefficients of image blocks, achieving an AUC above 0.95 for deepfake detection on datasets like FaceForensics++ [Giudice21]. More recent work combines noise residual autocorrelations with radial and angular spectrum averaging, achieving over 90% model attribution accuracy across GANs, VAEs, and diffusion models on diverse benchmarks including FFHQ and ImageNet subsets [Corvi23].
@@ II-B Existing Robustness Evaluations
While MFD methods perform well in benign settings, their robustness against adversarial evasion remains underexplored—particularly for model attribution, where attackers may attempt to remove or forge subtle, model-specific fingerprints. Existing evaluations largely examine resilience to generic image perturbations or targeted fingerprint removal, typically in black-box settings that do not adapt to the detector’s architecture. Moreover, most studies emphasize deepfake detection rather than attribution. Regarding the study of attacks, prior work has almost exclusively focused on fingerprint removal, overlooking forgery attacks in which adversaries mimic legitimate fingerprints to mislead attribution.
Generic Image Perturbations. Most evaluations test MFD robustness against generic image perturbations such as JPEG compression, blurring, resizing, and rotation, which can unintentionally degrade fingerprints. In the spatial pixel domain, Fourier spectrum–based methods have shown that JPEG compression can reduce detection accuracy by 10–20% on ProGAN datasets, primarily due to the loss of high-frequency components [Dzanic20]. DCT-based approaches demonstrate partial robustness, with accuracy declining by only 5–10% under scaling and rotation on FaceForensics++ (original AUC >> 0.95) [Giudice21]. CNN-based detectors are more resilient under moderate JPEG compression, retaining around 90% accuracy on social media-style datasets with CycleGAN images [Marra18]. Spectrum-domain perturbations, such as high-pass filters, have also been evaluated; for example, detectors like Frank et al.’s [Frank20] experienced up to a 30% accuracy drop on CelebA when subjected to such filters [Wesselkamp22]. Overall, black-box evaluations suggest that although some methods tolerate mild perturbations, others exhibit significant performance degradation even without adaptive adversarial strategies.
Method-Specific Attacks. A limited number of studies have investigated adaptive attacks specifically designed against MFD techniques, aiming to suppress fingerprints while maintaining perceptual quality. For co-occurrence-based detectors such as Nataraj et al.’s [Nataraj19], projected gradient descent (PGD) approximations of the co-occurrence operator can adjust pixel statistics to resemble those of real images, reducing detection accuracy from over 99% to nearly 0% in white-box settings on CycleGAN datasets [Goebel20]. Similarly, frequency-targeted attacks, including peak suppression and mean-spectrum subtraction, have been shown to evade detectors like Frank et al.’s [Frank20], lowering accuracy by 50–80% for deepfake detection on ProGAN and StyleGAN images [Wesselkamp22]. These results demonstrate the vulnerability of MFD methods to adaptive adversaries. However, the extent of their susceptibility to adaptive attacks, particularly against forgery and within attribution settings, remains largely underexplored.
II-C Research Gap
@@
    All prior research has focused on removal attacks, leaving robustness to forgery attacks underexplored. As we explain in Section III-B, although these two attacks are related, effectiveness of one does not imply the other.
@@ III-A System Model
This formulation—decoupling the fingerprint extractor $\phi$ from the attribution classifier hh—matches the structure of most MFD methods in the literature. Some approaches instead employ end-to-end neural networks, where fingerprint extraction is not a separate function (e.g., [Qian20, Wang20]). In such cases, the fingerprint can be defined implicitly: the preceding layers act as the fingerprint extractor $\phi$, while the final classification layer serves as the model attribution classifier $h$.
@@ III-B Adversarial Objectives
where $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ is a distance function and $\varepsilon$ specifies the perturbation budget. Typical choices of $d$ include $\ell_{\infty}$ or $\ell_{2}$ norms, LPIPS [zhang2018unreasonable], or PSNR. This constraint ensures that adversarial manipulations do not degrade the quality of the attacks images excessively.
@@
    A successful removal attack suppresses enough model-specific information in the fingerprint features to cause model attribution classifier $h$ to misclassify $x^{\prime}$. Since the MFD system operates in a closed world of candidate generators, such misclassification typically results in a low-confidence attribution to a different model.
@@ III-C Adversary Knowledge Levels
Note that removal attacks are feasible under all three knowledge levels, but forgery attacks are infeasible in the Black-box II setting because the adversary lacks access to $G_{t}$ and thus lacks the information needed to cause a targeted misattribution.
@@ IV-A White-box Attack Strategies
Figure 1 depicts which white-box attacks (W1–W3) apply given $\phi$’s properties. These strategies differ in their assumptions: W1 makes the strictest, requiring perfect gradient information. W2 and W3 are possible when there is gradient information but offer no advantage over W1. Similarly, W3 is possible when W2 is feasible, but not the other way around.
@@ IV-B Black-box Attack Strategies
  • B2: Image Transformations. Under Black-box Access II, without knowledge of $\phi$, $h$, or $\mathcal{Y}$, a typical adversary applies generic image transformations such as JPEG compression, noising, blurring, or resizing. These untargeted manipulations aim to degrade fragile fingerprint traces and reduce attribution confidence and without access to the target generator they cannot achieve forgery.
@@ IV-C Unified Optimization Objective
The adversarial loss $\mathcal{L}_{\text{adv}}$ is defined according attack objective:
@@ V-B MFD Methods
We evaluate 14 MFD methods spanning the RGB, frequency, and learned feature domains. Although this set is not exhaustive, it is representative: it includes influential methods that have been widely studied for model attribution (e.g., those covered in Song et al. [Song24]) and thus reflects the performance and robustness trends of the state of the art.
Table I summarizes their key characteristics, including the feature domain (RGB, frequency, or learned representations [Song24]), a brief description of the fingerprint features, and the fingerprint size (i.e., the output of $\phi$, or - for end-to-end methods). We also report attribution accuracy of the MFD methods, obtained by reproducing their evaluations on the new pool of 12 candidate generative models.
@@ V-C Models and Training Details
We train three neural network models to evaluate fingerprint robustness across different attack settings. All models are optimized with Adam for 100 epochs on $1,000$ images/fingerprints per generative model, using a batch size of 32, a validation split of 20%, and early stopping to prevent overfitting. The learning rate is fixed at $1\times 10^{-4}$.
@@
  • Attribution Classifier ($h$). This model maps fingerprint features to source model labels and serves as the primary target for attacks. It is implemented as a MLP with hidden dimensions [512, 256, 128] and ReLU activations, trained with cross-entropy loss. The architecture is fixed across all MFD methods, while the input dimensionality depends on the fingerprint size of the specific method.
@@ V-D Attack Implementation and Evaluation
We implement five adversarial attack strategies across the white-box (W1–W3) and black-box (B1–B2) threat models detailed in Section IV to evaluate fingerprint robustness. While removal attack are not targeted, forgery attacks require the attacker to induce a misclassification towards a specific target candidate generator label different from the true label. In the evaluation, we choose the target label uniformly at random on each execution of a forgery attack.
@@
PGD is limited to 50 iterations, optimizing the cross-entropy loss with a momentum coefficient of $0.9$, random initialization, and gradient clipping. The perturbation is bounded by $\epsilon=0.025$ (i.e., $\ell_{\infty}$ norm). We adopt the $\ell_{\infty}$ norm for its computational efficiency and its direct interpretation as a pixel value bound, while also verifying that perturbations remain visually acceptable. Across all attacks, this $\epsilon$ level consistently yields LPIPS <0.05<0.05 and PSNR >35>35 dB, which indicates satisfactory perceptual quality [zhang2018lpips, huynh2008psnr] (see Appendix A for more details). We also report the relationship between $\epsilon$ and ASR and LPIPS/PSNR in Figure 8 in Appendix D.
All gradient-based attacks support both removal and forgery objectives. For non-gradient attacks (B2), we employ four standard image perturbations with parameter settings chosen to preserve high image quality comparable to gradient-based attacks (i.e. LPIPS <0.05<0.05 and PSNR >35>35 dB). Specifically, we apply additive Gaussian noise with $\sigma=0.005$ (B2-Noise), Gaussian blur with $\sigma=0.5$ using a $3{\times}3$ kernel (B2-Blur), JPEG compression at quality $95$ (B2-JPEG), and isotropic downsampling to $0.9\times$ followed by upsampling (B2-Resize). Each perturbation is applied once independently to the original image, simulating one single adversarial query. We also conduct an ablation over these hyperparameters and report their impact on attack success in Figure 9 in Appendix D. Since B2’s perturbations are untargeted, we evaluate them only for removal attacks.
We conduct all experiments with PyTorch on NVIDIA A100 GPUs. The reported results are the mean and standard deviation over $5$ independent runs, where each run involves training all models in the pipeline and then conducting an attack. Upon acceptance, we will publish all our source code with documentation on how to reproduce the experiments.

### v2

@@ Abstract
Model fingerprint detection has shown promise to trace the provenance of AI-generated images in forensic applications. However, despite the inherent adversarial nature of these applications, existing evaluations rarely consider adversarial settings. We present the first systematic security evaluation of these techniques, formalizing threat models that encompass both white- and black-box access and two attack goals: fingerprint removal, which erases identifying traces to evade attribution, and fingerprint forgery, which seeks to cause misattribution to a target model. We implement five attack strategies and evaluate 14 representative fingerprinting methods across RGB, frequency, and learned-feature domains on 12 state-of-the-art image generators. Our experiments reveal a pronounced gap between clean and adversarial performance. Removal attacks are highly effective, often achieving success rates above $80\%$ in white-box settings and over $50\%$ under black-box access. While forgery is more challenging than removal, its success varies significantly across targeted models. We also observe a utility–robustness trade-off: accurate attribution methods are often vulnerable to attacks and, although some techniques are robust in specific settings, none achieves robustness and accuracy across all evaluated threat models. These findings highlight the need for techniques that balance robustness and accuracy, and we identify the most promising approaches toward this goal. Code available at: https://github.com/kaikaiyao/SmudgedFingerprints.
@@ I Introduction
The widespread adoption of generative AI models for image generation has created an urgent need for content provenance mechanisms. These models—now integrated into creative tools [adobe2023firefly], content platforms [katsamakas2024gcp], and commercial APIs [openai2023dalle3]—can produce images that are almost visually indistinguishable from authentic photographs. While these technological advances enable new forms of creativity, they also come with risks for misinformation, fraud, and abuse, particularly when AI-generated content is passed as genuine [bird2023typology]. As a consequence, reliably identifying and tracing generated images has become a pressing challenge for maintaining trust and accountability in digital media.
@@
Existing MFD techniques evaluated under various settings achieve over 90% accuracy for both deepfake detection [Marra19a, Nataraj19, Durall20, Dzanic20, Wang20, Corvi23] and model attribution [Marra18, Marra19b, Yu19, Wang20, Guarnera20, Girish21]. If these accuracies were to hold in realistic scenarios, MFD would prove useful in bridging the accountability gap in current deployments of generative models. Indeed, being able to identify deepfakes would help mitigate the risks of misinformation and scams, while attributing a harmful output to the specific model that generated it would make it possible to flag issues, understand their causes, and prevent similar issues in the future.
@@
We focus on model attribution rather than deepfake detection because model attribution is the harder task for MFD: it requires not merely detecting the presence of model-induced patterns, but distinguishing architecture-specific differences across them, which requires higher entropy. Therefore, evaluating MFD techniques for model attribution allows us to draw conclusions for deepfake detection: if the techniques prove robust against attacks on model attribution, they should remain robust in detecting for the simpler task of deepfake detection.
@@ II-A Model Fingerprint Detection
Model fingerprint detection (MFD) has gained significant attention as a passive approach to identifying images produced by generative models, such as GANs, VAEs, and diffusion models. Unlike watermarking, which embeds signals during generation, MFD relies on detecting inherent artifacts or “fingerprints” left by the model’s processes. These fingerprints can be detected in statistical irregularities of color and frequency patterns in the image.
@@
Frequency Domain. Generative models often introduce spectral anomalies, such as deficiencies in high-frequency components or unnatural power distributions, which frequency-based methods target through transformations like FFT or DCT. For example, approaches using azimuthally averaged power spectral densities from 2D FFTs have demonstrated near-perfect discrimination (close to 100% accuracy) for deepfake detection between GAN outputs and real images on CelebA and LSUN, detecting patterns in high-frequency components for models like DCGAN and CycleGAN [Durall20]. Other methods extract features from Fourier modes, such as decay exponents and magnitudes, achieving up to 99% accuracy for deepfake detection on ProGAN-generated faces [Dzanic20]. Giudice et al. instead rely on DCT analysis, extracting a 63-dimensional fingerprint vector from the means of AC coefficients of image blocks, achieving an AUC above 0.95 for deepfake detection on datasets like FaceForensics++ [Giudice21]. More recent work combines noise residual autocorrelations with radial and angular spectrum averaging, achieving over 90% model attribution accuracy across GANs, VAEs, and diffusion models on diverse benchmarks including FFHQ and ImageNet subsets [Corvi23].
@@ II-B Existing Robustness Evaluations
While MFD methods perform well in benign settings, their robustness against adversarial evasion remains underexplored—particularly for model attribution, where attackers may attempt to remove or forge subtle, model-specific fingerprints. Existing evaluations mostly test robustness against generic image perturbations or targeted fingerprint removal, typically in black-box settings that do not adapt to the detector’s architecture. Moreover, most studies emphasize deepfake detection rather than model attribution. Regarding the study of attacks, prior work has almost exclusively focused on fingerprint removal, overlooking forgery attacks in which adversaries mimic legitimate fingerprints to mislead attribution.
Generic Image Perturbations. Most evaluations evalate MFD robustness to generic image perturbations such as JPEG compression, blurring, resizing, and rotation, which can unintentionally degrade fingerprints. In the spatial pixel domain, Fourier spectrum–based methods have shown that JPEG compression can reduce detection accuracy by 10–20% on ProGAN datasets, primarily due to the loss of high-frequency components [Dzanic20]. DCT-based approaches demonstrate partial robustness, with accuracy declining by only 5–10% under scaling and rotation on FaceForensics++ (original AUC >> 0.95) [Giudice21]. CNN-based detectors are more resilient under moderate JPEG compression, retaining around 90% accuracy on social media-style datasets with CycleGAN images [Marra18]. Spectrum-domain perturbations, such as high-pass filters, have also been evaluated; for example, detectors like Frank et al.’s [Frank20] experienced up to a 30% accuracy drop on CelebA when subjected to such filters [Wesselkamp22]. Overall, black-box evaluations suggest that although some methods tolerate mild perturbations, others exhibit significant performance degradation even without adaptive adversarial strategies.
Method-Specific Attacks. A few studies have investigated adaptive attacks specifically designed against MFD techniques, aiming to suppress fingerprints while maintaining perceptual quality. For co-occurrence-based detectors such as Nataraj et al.’s [Nataraj19], projected gradient descent (PGD) approximations of the co-occurrence operator can adjust pixel statistics to resemble those of real images, reducing detection accuracy from over 99% to nearly 0% in white-box settings on CycleGAN datasets [Goebel20]. Similarly, frequency-targeted attacks, including peak suppression and mean-spectrum subtraction, have been shown to evade detectors like Frank et al.’s [Frank20], lowering accuracy by 50–80% for deepfake detection on ProGAN and StyleGAN images [Wesselkamp22]. These results demonstrate the vulnerability of MFD methods to adaptive adversaries. However, the extent of their susceptibility to adaptive attacks, particularly against forgery and within model attribution settings, remains largely underexplored.
II-C Research Gaps
@@
    All prior research has focused on removal attacks, leaving robustness to forgery attacks underexplored. As we explain in Section III-B, although these two attacks are related, success of one does not imply the other.
@@ III-A System Model
This formulation—decoupling the fingerprint extractor $\phi$ from the attribution classifier hh—matches the structure of most MFD methods in the literature. Some approaches instead employ end-to-end neural networks, where fingerprint extraction is not a separate function (e.g., [Qian20, Wang20]). In such cases, the fingerprint can be defined implicitly: the neural network except the last layer acts as the fingerprint extractor $\phi$, and the last layer can be interpreted as the model attribution classifier $h$.
@@ III-B Adversarial Objectives
where $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ is a distance function and $\varepsilon$ specifies the perturbation budget. Typical choices of $d$ include $\ell_{\infty}$ and $\ell_{2}$ norms, LPIPS [zhang2018unreasonable], and PSNR. This constraint ensures that adversarial manipulations do not degrade the quality of the attacked images excessively.
@@
    A successful removal attack suppresses enough model-specific information in the fingerprint features to cause the model attribution classifier $h$ to misclassify $x^{\prime}$. Since the MFD system operates in a closed world of candidate generators, such misclassification typically results in a low-confidence attribution to a different model.
@@ III-C Adversary Knowledge Levels
Removal attacks are feasible under all three knowledge levels, while forgery attacks are infeasible in the Black-box II setting, as the adversary does not have access to $G_{t}$, and thus lacks the information needed to cause a targeted misattribution.
@@ IV-A White-box Attack Strategies
Figure 1 depicts which white-box attacks (W1–W3) apply given properties of $\phi$. These strategies differ in their assumptions: W1 is the strictest, requiring perfect gradient information. W2 and W3 are possible when there is gradient information but offer no advantage over W1. Similarly, W3 is possible when W2 is feasible, but not the other way around.
@@ IV-B Black-box Attack Strategies
  • B2: Image Transformations. Under Black-box Access II, without knowledge of $\phi$, $h$, or $\mathcal{Y}$, a typical adversary applies generic image transformations such as JPEG compression, noising, blurring, or resizing. These untargeted manipulations aim to degrade fragile fingerprint traces and reduce attribution confidence.
@@ IV-C Unified Optimization Objective
The definition of the adversarial loss $\mathcal{L}_{\text{adv}}$ depends on the attack objective:
@@ V-B MFD Methods
We evaluate 14 MFD methods spanning the RGB, frequency, and learned feature domains. Although this set is not exhaustive, it is representative: it includes influential MFD methods that have been widely studied for model attribution (e.g., those covered in Song et al. [Song24]) and thus reflects the performance and robustness trends of the state of the art.
Table I summarizes their key characteristics, including the feature domain (RGB, frequency, or learned representations [Song24]), a brief description of the fingerprint features, and the fingerprint size (i.e., the output of $\phi$, or - for end-to-end methods). We also report attribution accuracy of the MFD methods, obtained by reproducing their evaluations on the pool of 12 candidate generative models.
@@ V-C Models and Training Details
We train three neural network models to evaluate fingerprint robustness across different attack settings. All models are optimized with Adam for 100 epochs on 1,0001{,}000 images/fingerprints per generative model, using a batch size of 32, a validation split of 20%, and early stopping to prevent overfitting. The learning rate is fixed at $1\times 10^{-4}$.
@@
  • Attribution Classifier ($h$). This model maps fingerprint features to source model labels and serves as the primary target for attacks. It is implemented as an MLP with [512, 256, 128] hidden dimensions and ReLU activations, trained with cross-entropy loss. The architecture is fixed across all MFD methods, while the input dimensionality depends on the fingerprint size of the specific method.
@@ V-D Attack Implementation and Evaluation
We implement five adversarial attack strategies across the white-box (W1–W3) and black-box (B1–B2) threat models detailed in Section IV to evaluate fingerprint robustness. While removal attacks are not targeted, forgery attacks require the attacker to induce a misclassification towards a specific target candidate generator label different from the true label. In the evaluation, we choose the target label uniformly at random on each execution of a forgery attack.
@@
PGD is limited to 50 iterations, optimizing the cross-entropy loss with a momentum coefficient of $0.9$, random initialization, and gradient clipping. The perturbation is bounded by $\epsilon=0.025$ (i.e., $\ell_{\infty}$ norm). We adopt the $\ell_{\infty}$ norm for its computational efficiency and its direct interpretation as a pixel value bound, while allowing to verify that perturbations remain visually acceptable. Across all attacks, this $\epsilon$ level consistently yields LPIPS <0.05<0.05 and PSNR >35>35 dB, which indicates satisfactory perceptual quality [zhang2018lpips, huynh2008psnr] (see Appendix A for more details). We also report the relationship between $\epsilon$ and ASR and LPIPS/PSNR in Figure 8 in Appendix D.
All gradient-based attacks support both removal and forgery objectives. For non-gradient attacks (B2), we employ four standard image perturbations with parameter settings chosen to preserve high image quality comparable to gradient-based attacks (i.e., LPIPS <0.05<0.05 and PSNR >35>35 dB). Specifically, we apply additive Gaussian noise with $\sigma=0.005$ (B2-Noise), Gaussian blur with $\sigma=0.5$ using a $3{\times}3$ kernel (B2-Blur), JPEG compression at quality $95$ (B2-JPEG), and isotropic downsampling to $0.9\times$ followed by upsampling (B2-Resize). Each perturbation is applied once independently to the original image, each simulating one single removal attempt. We also conduct an ablation over these hyperparameters and report their impact on attack success in Figure 9 in Appendix D. Since B2’s perturbations are untargeted, we evaluate them only for removal attacks.
We conduct all experiments with PyTorch on NVIDIA A100 GPUs. The reported results are the mean and standard deviation over $5$ independent runs, where each run involves training all models in the pipeline and then conducting an attack.